How many milliliters of a 0.266 M LiNO3 solution are required to make 150.0 mL of 0.075 M LiNO3 solution?

42.3 mL

53.2 mL

18.8 mL

23.6 mL

35.1 mL
Answer #1

The prepared solution have 0.075M of LiNO3

As we know the molarity is the number of moles of solute present in the 1L of solution.

molarity=\frac{w}{m}\times \frac{1000}{v}

Here the prepared solution have

w= mass= to be got

m= molar mass of LiNO3= 68.94g/mol

v=volume = 150mL

Molarity= 0,075M

By applying the law molarity=\frac{w}{m}\times \frac{1000}{v}

0.075M=\frac{w}{68.94}\times \frac{1000}{150}\\ w= \frac{150\times 68.94\times 0.075}{1000}\\ w= \frac{775.575}{1000}\\ w=0.775g

Hence the prepared solution has 0.775g of LiNO3solute.

Now for the preparation of 150mL of 0.075M LiNO3 we need 0.266M of LiNO3 that carry 0.775g of LiNO3. For this, the required volume of the solution is x mL(suppose).

We need to find out the value of x

Here the given data are

volume (v)= x mL

Molarity= 0.266M

Mass(w)= 0.775g

Molar mass of LiNO3= 68.94g/mol

Now apply the rule molarity=\frac{w}{m}\times \frac{1000}{v}

0.266M=\frac{0.775}{68.94}\times \frac{1000}{x}\\ 0.266M=0.0112\times \frac{1000}{x}\\ 0.266M\times x=0.0112\times1000\\ 0.266M\times x=11.24\\ x=\frac{11.24166}{0.266}\\ x=42.3mL

Hence the volume is 42.3 mL.

So by seeing the above expression the correct answer is 42.3 mL.